One path on a project has 3 tasks, A, B, and C. The expected times for these tasks are 31, 34, and 20 respectively. Their standard deviations are 3, 4 and 2 respectively. The expected time for this path is                             [ Select ]                       ["65", "95", "75", "85"]       and standard deviation of this path is                             [ Select ]                       ["9", "9.7", "5.4", "29"]         .
Answer #1

The expected time for the path = Sum of each Activity = ( 31 + 34 + 20) = 85

Hence, the correct option is 85.

Stabdard deviation = Sum of all standard deviation = 3 + 4 + 2 = 9.

Hence, the correct option is 9.

Do give a positive rating if you think this answer helped